# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

### v2

@@ Abstract
In this work, we propose a unified representation for Super-Resolution (SR) and Image Compression, termed Factorized Fields, motivated by the shared principles between these two tasks. Both SISR and Image Compression require recovering and preserving fine image details—whether by enhancing resolution or reconstructing compressed data. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition to explicitly capture multi-scale visual features and structural components in images, addressing the core challenges of both tasks. We first derive our SR model, which includes a Coefficient Backbone and Basis Swin Transformer for generalizable Factorized Fields. Then, to further unify these two tasks, we leverage the strong information-recovery capabilities of the trained SR modules as priors in the compression pipeline, improving both compression efficiency and detail reconstruction. Additionally, we introduce a merged-basis compression branch that consolidates shared structures, further optimizing the compression process. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA.
@@ 1 Introduction
Figure 1: The Proposed Factorized Fields. With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, especially useful for both Super-Resolution and Image Compression. Note that in this figure $\alpha=\{1,16\}$ and $\psi=sin$.
@@
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches [46, 21]laid the foundation for SISR, which was later enhanced with GAN-based methods [39] for improved perceptual realism. Follow-up Transformer-based networks [12] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [19, 15, 16], which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [33, 69, 66] lay the groundwork. The emerging learned image compression models [7, 2, 29, 48, 58, 17], compression algorithms of which mostly follow the pixel-space transform coding  [7, 27]paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
@@
Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [9, 59, 8, 24, 6, 10, 26], we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation. In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements. On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
Finally, we propose merging the bases by introducing an additional compression branch, consolidating multiple bases into one alongside multi-image transmission. This approach leverages the mutual information across multiple images, reducing the need for redundant transmissions and refining the basis structure. The main contributions of this paper are summarized as follows:
  • We propose Generalized Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
  • We integrate super-resolution with image compression by introducing SR prior during decompression to compensate for lost details and developing a merged-basis compression branch for multi-image compression.
@@ 2 Related Works
Super-Resolution (SR).
Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs. Following the foundational studies, CNN-based strategies [21, 46, 79, 64, 34] were initially introduced with modeling techniques such as residual learning [40, 47, 68, 78, 77, 18, 45] , or recursive learning [16, 65]. Besides, subsequent research also sheds light on GAN-base methods [39, 40, 70, 71] to enhance realisticity and detail quality. However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks [12, 43]. Afterward, SwinIR [44] is proposed to combine spatial locality and non-local information by Swin Transformer [50] with window attention and achieve breakthrough improvement in SISR. Following SwinIR’s success, several works have built upon its framework [19, 80, 75] to reach better image quality as well as solve information bottleneck [18]. Hybrid approaches CRAFT [41] merge the benefits of convolutional and transformer structures to further elevate SR performance. For better feature aggregation, DAT [15] and HAT [13] integrate spatial and channel information using attention mechanisms to enhance their representation capabilities. Moreover, RGT [16] introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
Image Compression (IC).
Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG [69] and JPEG-2000 [66]. Early CNN-based approaches [7, 3, 28] have been surpassed by transformer-based models [37, 81, 72, 22], which leverage spatio-channel attention for better performance. GAN-based methods [53, 54, 23, 62] have further contributed to real-time adaptive compression. Recently, ELIC [30] introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM [48] integrated transformers and CNNs to capture both global and local image features. The eContextformer [38] introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism. GroupedMixer [42] proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling. Meanwhile, Wavelet Conditional Diffusion [63] introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
In summary, while recent super-resolution and image compression advances focus on increasingly complex architectures, our work takes a different approach. We propose Factorized Fields, a unified framework that models visual and structural features, offering a more comprehensive solution to enhance performance in both tasks.
@@ 3 Methods
In this section, we present the details of our proposed methods. Inspired by factor fields, which can be formulated by
and whose details can be found in supplementary materials , we first explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in  Sec. 3.1. We then describe how to adapt the formulation to Super-Resolution in  Sec. 3.2. Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in  Sec. 3.3. Note that the background of learned image compression is described in supplementary materials.
@@
Figure 3: The overall pipeline of image super-resolution with our Factorized Fields. Given a low-resolution image $I_{LR}$, we first extract coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficient and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 4, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
3.1 Formulation of Factorized Fields
As discussed in  Sec. 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details. Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information [25, 36] to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation. Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
Inspired from recent success on such decomposition fields [24, 10, 9, 59, 6, 8], we primarily built our Factorized Fields framework on factor fields [10] from  Eq. 1:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
Note that $\hat{I}$ denotes the approximated images, and $x\in R^{2}$ are the pixel coordinates.
Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated.
A key insight of our work is the use of sawtooth coordinate transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, which performs best among tested transformations in image regression tasks. The sawtooth transformation maps input coordinates to a periodic space, implicitly capturing patch-like frequency information as shown in Fig. 2a. However, traditional downsampling in neural networks fails to preserve this periodic structure. To address this, we introduce sawtooth-aware downsampling as in Fig. 2c with our Basis Swin Transformer. This rearrangement allows the network to learn correlations between the periodic features in Sawtooth sampling.
Moreover, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images. To resolve this, we propose a modified version, our Factorized Fields:
@@ 3.1 Formulation of Factorized Fields
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (3)
where $\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R$. Here, inspired by the Fourier Series, with a scalar $\alpha$ which corresponds to an increase in sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
Refer to captionRefer to caption
Figure 4: Comparison of Different Representation This figure illustrates the performance comparison on a single-image regression task using three different representations: vanilla factor fields in Eq. 1, our Factorized Fields (Eq. 3), and Plain Fields, which use simple channel-wise tensor concatenation instead of coefficient-basis multiplication. The left panel shows the MAE on the magnitude of different frequency components obtained from the Fourier Transform, while the right panel shows PSNR along training. Although the basis-coefficient decomposition method in [10] yields improvements, it still struggles with slow convergence and inconsistencies between low and high-frequency components. In contrast, by introducing explicit formulations for $\alpha$ and $\psi$, our method aligns different frequency components effectively and achieves superior convergence speed. This advantage becomes especially crucial when the task extends to generalizable applications, such as super-resolution or image compression.
Our Factorized Fields enhance the original Factor Fields [10] in three key ways: (1) non-linear transformation through $\psi\in\{\text{sin},\text{cos}\}$ enables better modeling of complex image patterns, (2) multi-frequency modulation using $\alpha_{j}$ allows explicit control over frequency components, and (3) sawtooth-aware downsampling preserves periodic structure during feature processing. As shown in Fig. 4, our refined representation contributes to large margin of improvement in convergence and performance, demonstrating the modification is the key to superior information modeling.
@@
Figure 5: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec. 3.2 as information-recovery prior, detailed in Sec. 3.3.1. (c) Expanding on (b),  Sec. 3.3.2 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
3.2 Super-Resolution with Factorized Fields

### v3

# FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models

@@ Abstract
In this work, we propose using a unified representation, termed Factorized Features, for low-level vision tasks, where we test on Single Image Super-Resolution (SISR) and Image Compression. Motivated by the shared principles between these tasks, they require recovering and preserving fine image details—whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition as well as an explicit formulation of frequencies to capture structural components and multi-scale visual features in images, which addresses the core challenges of both tasks. We replace the representation of prior models from simple feature maps with Factorized Features to validate the potential for broad generalizability. In addition, we further optimize the pipelines by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression and super-resolution. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA. Please see our project page at jayisaking.github.io/FIPER.
@@ 1 Introduction
Figure 1: The Proposed Factorized Features. With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression. Note that in this figure $\alpha=\{1,16\}$ and $\psi=\text{sin}(\cdot)$.
@@
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches [52, 23] laid the foundation for SISR, which was later enhanced with GAN-based methods [40] for improved perceptual realism. Follow-up Transformer-based networks [14] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [21, 17, 18], which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR? On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [35, 82, 79] lay the groundwork. The emerging learned image compression models [8, 3, 30, 55, 67, 19], compression algorithms of which mostly follow the pixel-space transform coding  [8, 28] paradigm, then introduce a variety of networks  [83, 55] or difference losses [31] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
@@
Based on the aforementioned analysis, it becomes clear that although these tasks appear to be distinct, they share mutual similarities in two key aspects: (1) The tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) They aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [11, 68, 9, 26, 7, 12, 27], we propose a unified representation, Factorized Features, with generalizable Coefficient Backbone and Basis Transformer for learned coefficient and basis, respectively. This approach explicitly captures multi-scale visual features and repetitive structral components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation.
Finally, we propose merging the bases for multi-image processing by consolidating multiple bases into one through a network, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
The main contributions of this paper are summarized as follows:
  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in Sec. 3.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
  • We generalize such image descriptor to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
@@ 2 Related Works
Super-Resolution (SR). Image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, which is crucial for computer vision. Early CNN-based methods [23, 52, 95, 76, 36] introduced residual [41, 54, 81, 94, 93, 20, 51] and recursive learning [18, 77]. GAN-based approaches [40, 41, 84, 85] improved perceptual quality but faced spatial locality constraints, leading to Transformer-based SISR models [14, 46]. SwinIR [48] integrated window attention with Swin Transformer [58], inspiring further advancements [21, 96, 91, 20]. Hybrid models like CRAFT [43], DAT [17], and HAT [15] optimize feature aggregation, while RGT [18] enhances spatial details efficiently.
Image Compression (IC). Deep learning surpasses traditional compression methods like JPEG [82] and JPEG-2000 [79], with CNN-based [8, 4, 29], Transformer-based [38, 97, 86, 24], and GAN-based [61, 62, 25, 73] approaches improving performance. ELIC [32] introduced adaptive coding, LIC TCM [55] combined CNNs and Transformers, and eContextformer [39] leveraged spatial-channel attention. GroupedMixer [44] introduced token mixers, while Wavelet Conditional Diffusion [75] balanced perceptual quality and distortion.
Our work diverges from these trends by introducing Factorized Features, a unified framework modeling both visual and structural features to enhance performance across super-resolution and image compression.
@@ 3 Methods
In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [57], and Factor Fields [12], and delve into the motivation and how we resolve the weaknesses of these formulations by our Factorized Features in Sec. 3.1. Next, we discuss how to integrate such strategy into learned settings of Super-Resolution in Sec. 3.2 and Image Compression Sec. 3.3. Note that the background of learned image compression is provided in supplementary materials.
@@
Figure 3: The overall pipeline of image super-resolution with our Factorized Features. Given a low-resolution image $I_{LR}$, we first extract the coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficients and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, and $W$ denote the scale factor, height, and width, respectively.
3.1 Formulation of Factorized Features
Fourier Transform has been long used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions. In practice, it often takes the form of Discrete Cosine Transform for real-valued 2D discrete signal $f$ with height $H$ and width $W$:
$f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right]$ (1)
where $S_{u,v}$ is the amplitude.
Without the loss of generality, we consider square images in this work, with edge length (and also the period) denoted as $T$:
$f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right]$ (2)
where $x<T$, $y<T$, and $\xi_{u,v}$ is phase shift. In image processing tasks, the goal is to learn $S$ and $\xi$, then use the formula above to reconstruct high-quality images from degraded inputs. To achieve this, [57] treats $S$ and $\xi$ as learnable parameters, allowing the model to optimize them for better reconstruction. Furthermore, the frequencies $u$ and $v$ are also designed to be promptable by the model, enabling dynamic adaptation to spatial-varying high- and low-frequency components.
More generally, we can formulate it as the multiplication of coefficient $c$ and basis $b$ with simplified notation, and in this work, we use $\mathbf{x}$ to represent the pixel coordinates $(x,y)$:
$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),$ (3)
where $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ and $N$ denotes the number of frequency components.
Spatially Variant Coefficient. Intuitively, the local frequency spectrum of a signal is spatially variant, meaning that different regions exhibit distinct spectral characteristics, e.g., certain regions of an image do not require high-frequency components. Therefore, in this work, we argue that coefficient $c_{i}$ should be spatial-varying, which is different from previous works, i.e.,
$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
Learned Non-uniform Basis. Previous works learn basis through frequencies ${u,v}$ and phase shift $\xi$[57, 47, 42]; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details; in other words, since low-frequency components typically require more energy [45], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details, and the predefined nature of these sinusoidal functions further restricts local flexibility, making fine-grained adjustments more challenging. As shown in Fig. 5, higher frequency components converge slower than their counterpart. Hence, in this work, we generate the entire basis map $b_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map and combine them together. In this work, the composition process happens in the end with a projection function $\mathcal{P}$ to form RGB images $\hat{I}$:
$\hat{I}(\mathbf{x})=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl{\}}\Bigl{)}.$ (5)
In contrast to previous works which employ fourier bottlenecks[57], or frequency decomposition blocks[47, 42, 45] among model blocks, here we seek to address a fundamental question: Can a well-learned representation effectively replace the simple feature map output ? That is, our solution is meant to be applied directly on networks to represent any images.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Eq. 4 (b) Eq. 6 (c) Eq. 7 with only $\psi$ (d) Eq. 7 with both $\psi$ and $\alpha$
Figure 4: Visualization of different factorization methods.
Coordinate Transformation. The success of Fourier-based representation originates from its sinusoidal decomposition. By using such periodic functions, one can effectively analyze and reconstruct signals with the enforcement of high- and low-frequency components. However, the formulation in Eq. 5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in Fig. 4 we witness performance loss due to lack of such guidance. Inspired by position encoding [65, 11, 68, 7, 9] and Fourier feature mapping [78], we introduce coordinate transformation function $\gamma_{i}$ on $\mathbf{x}$ before sampling basis $b_{i}$:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, visualized in Fig. 2.a. We argue that the function explicitly imposes patch-like periodic constraints and, therefore, enforces models to learn such repetitive patterns.
Multi-frequency Modulation. [72] has shown that neural networks are biased towards learning low-frequency content. Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in Fig. 4:
@@ 3.1 Formulation of Factorized Features
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (7)
where $\psi\in{\sin,\cos}$ and $\alpha\in R$. In this way, each basis $b_{i}$ contributes to both high- and low-frequency components by different values of $\alpha$. Consequently, if a basis is only accurate in the low-frequency domain, the output will contain undesired high-frequency noise.
Alternatively, Eq. 7 can be interpreted as making the spatial coordinates $x,y$ in Eq. 1 learnable instead of the frequency components $u,v$. To validate this, we conduct experiments (Fig. 5) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
To this end, our full Factorized Features representation is formulated in Eq. 7.
Figure 5: In this figure, we validate the claim that fixed sinusoidal functions attend more to low-frequency parts. Learnable basis denotes the Learned Non-uniform basis in Sec. 3.1. And we can see that Learnable Coordinate $x$ has the same effect, verifying our conclusion.
@@
Figure 6: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec. 3.2 as information-recovery prior, detailed in Sec. 3.3 (c) Expanding on (b),  Sec. 3.3.1 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
3.2 Super-Resolution with Factorized Features